A coupon bond pays $100 at the end of each year for 3 years, and at the maturity date in 3 years, in addition to the final coupon payment, the bond also makes a face value payout of $500. If the interest rate is 3% then this bond has a present value of $740.43 $755.55 $762.98 $808.08
Answer #1

Answer:Option A:$740.43

PV of Bond=C/(1+r)^1+C/(1+r)^2+C/(1+r)^3+P/(1+r)^3

C= Coupon rate

P=face value

r=interest rate

n=years to maturity

PV of Bond=100/(1+3%)^1+100/(1+3%)^2+100/(1+3%)^3+500/(1+3%)^3=$740.43